Flying to Kampala with a tailwind a plane averaged 158 km/h. On the return trip the plane only averaged 112 km/h while flying back into the same wind. Find the speed of the wind and the speed of the plane in still air.

Answer by mananth(16946)

x= Plane speed in still air
y= wind speed

1 x + 1 y = 158 .............1
Total value
1 x -1 y = 112 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
1 x y = 158
x - y = 112
Add the two equations
2 x = 270
/ 2
x = 135
plug value of x in (1)
1 x + 1 y = 158
135 + y = 158
y = 158 -135
y = 23
y = 23
x= 135 Plane speed in still air 135
y= 23 wind speed 23
m.ananth@hotmail.ca